COMPLETELY REGULAR MULTIVARIATE STATIONARY PROCESS AND THE MUCKENHOUPT CONDITION

SERGEI TREIL ALEXANDER VOLBERG

Abstract. We give necessary and sufficient conditions for a multivariate stationary stochastic process to be completely regular. We also give the answer to a question of V.V. Peller concerning the spectral measure characterization of such processes.